Please help me: the triangles hypotenuse is r and the other short sides are x and y. solve for the third side of the triangle if x= 2 sqrt ( 5 ) and r= sqrt ( 38 ). Thanks!
Answer by stanbon(75887)

the triangles hypotenuse is r and the other short sides are x and y. solve for the third side of the triangle if x= 2 sqrt ( 5 ) and r= sqrt ( 38 ).
----------
Using Pythagoras:
r^2 = x^2 + y^2
38 = 4*5 + y^2
y^2 = -7
---
That would make y a complex number.
---
Conclusion: No Real Number solution.
===================================
Cheers,
Stan H.